Compute the amount to be paid for each of the four separate invoices assuming that all invoices are paid within the discount period.

Merchandise (Gross) Terms

a. $9200 2/10, n/60

b. $23600 1/15, EOM

c. $83400 1/10, n/30

d. $20500 3/15, n/45
Answer #1

Assuming all invoices are paid within the discounting period, calculation of amount to be paid for each invoice is as follows:
a) $9200 * (1 - 0.02) = $9,016 i.e. after 2% discount when paid within 10 days.
b) $23600 * (1 - 0.01) = $23,364 i.e. after 1% discount when paid within 15 days.
c) $83400 * (1 - 0.01) = $82,566 i.e. after 1% discount when paid within 10 days.
d) $20500 * (1 - 0.03) = $19,885 i.e. after 3% discount when paid within 15 days.